If x = t^3 + y^3 where t= a cos v , y= Vsin v. Find dx/dv


1
Expert's answer
2021-11-23T09:21:58-0500
"dx\/dv=dx\/dt(dt\/dv)+dx\/dy(dy\/dv)"

"=3t^2\big(-a\sin (v)\big)+3y^2\big(\sin (v)+v\cos(v)\big)"

"=-3a\big(a\cos(v)\big)^2\sin(v)"

"+3\big((v\sin(v)\big)^2\big(\sin(v)+v\cos(v)\big)"

"=-3a^3\sin(v)\cos^2(v)+3v^2\sin^2(v)\big(\sin(v)+v\cos(v)\big)"

Or


"x(v)=a^3\cos^3(v)+v^3\sin^3(v)"


"dx/dv=-3a^3\sin(v)+\cos^2(v)"

"+3v^2\sin^3(v)+3v^3\sin^2(v)\cos(v)"